Problem 7:

The Fed conducts $20 million open-market purchase of government bonds. If the required reserve ratio is 10 percent, what is the largest possible increase in the money supply that could result? Explain. What is the smallest possible increase? Explain.
Answer #1

When Reserve requirement is 10%

The multiplier = 1/R = 1/0.10 = 10

The open market purchase of 20 million will increase the money supply by

= 20*10

= 200 million